# A minimal Gröbner basis for simple $\mathfrak{sl}_{n}$- or $\mathfrak{sp}_{n}$-modules

### 03 Jan 2024 v1

@@ Abstract.
We explicitly provide minimal Gröbner bases for simple, finite-dimensional modules of complex Lie algebras of types A and C, using a weighted ordering that is compatible with the PBW filtration on the universal enveloping algebras.
@@ 1. Introduction
as an element of $M_{\lambda}$. Since $f_{\alpha_{2}}^{2}$ is also contained, the set is not reduced. Paragraph 1: ”These computations are carried out using OSCAR [Osc] and the implementation of the Gröbner basis for other types can be found in the latter.
@@
Acknowledgments The first author gratefully acknowledges financial support by the DFG –Project-ID 286237555–TRR 195.
@@ 2.3. Gröbner bases
For the following, we need to introduce a bit of notation: Let $\lambda=\sum m_{i}\omega_{i}\in P^{+}$, we set $\lambda_{i}=\sum_{j=i}^{n}m_{j}$. Then $\lambda$ is also the partition $(\lambda_{1}\geq\lambda_{2}\geq\ldots\geq\lambda_{n}\geq 0)$. Let $\mathbf{s}\in\mathbb{R}^{|R^{+}|}$ be supported on the roots of the subdiagram induced from $\alpha_{\ell},\ldots,\alpha_{k}$. For type $A$ we set
@@
$\partial_{\mathbf{s}}:=e_{\ell,\ell}^{s_{\ell+1,\bullet}}\cdots e_{\ell,k-1}^{s_{k,\bullet}}e_{k,k}^{s_{\bullet,k-1}}\cdots e_{\ell+1,k}^{s_{\bullet,\ell}}\in U(\mathfrak{n}^{+})$
To ease the notation, we denote $\partial_{\mathbf{s}}$ the similar element in type $C$ as defined in [FFoL11a, Theorem 3.4]. We set
$d(\mathbf{s})=(s_{n,\bullet},\ldots,s_{1,\bullet})\in\mathbb{Z}^{n}$
and $>$ to be the lexicographic ordering on $\mathbb{Z}^{n}$.
Since $I_{\lambda}$ is finitely generated and $U(\mathfrak{g})$ is a G-algebra (also known as PBW algebra or algebra of solvable type), a (left-)Gröbner basis exists for each monomial ordering (see [KRW90]). We recall the monomial ordering on $U(\mathfrak{n}^{-})$ introduced in [FFoL11, FFoL11a] here, let us briefly write $f_{i,j}$ for $f_{\alpha_{i,j}}$:
$\begin{array}[]{rcl}f_{n,n}&>&\\ f_{n-1,\overline{n-1}}>f_{n-1,n}&>&f_{n-1,n-1}>\\ \ldots&\ldots&\ldots\\ f_{1,\overline{1}}>f_{1,\overline{2}}>\ldots>f_{1,n}&>&f_{1,n-1}>\ldots>f_{1,1}\end{array}$
We use the same notation for the induced homogeneous lexicographic ordering on $U(\mathfrak{n}^{-})$. We will use $>$ to introduce another monomial ordering $\succ$ ([FFoL11a, Remark 2.7]): For two monomials $f^{\mathbf{s}},f^{\mathbf{t}}$ we say $f^{\mathbf{s}}\succ f^{\mathbf{t}}$ if either $|\mathbf{s}|>|\mathbf{t}|$ or if they have the same total degree $d(\mathbf{s})<d(\mathbf{t})$ or if they are also equal $f^{\mathbf{s}}>f^{\mathbf{t}}$.
We extend this to $\mathfrak{g}$ such that $\mathfrak{n}^{-}\succ\mathfrak{h}\succ\mathfrak{n}^{+}$. We consider the induced weighted ordering $\geq$ by setting degree of $\mathfrak{n}^{-}$ equals $1$ and degree of $\mathfrak{n}^{+}\oplus\mathfrak{h}$ equals 00.
Let $\mathbf{s}$ be supported on a Dyck path $\mathbf{p}\in D_{i,}$, then it is shown in [FFoL11, FFoL11a], that $LM(\partial_{\mathbf{s}}f_{i,j}^{\text{deg }\mathbf{s}})=f^{\mathbf{s}}$.
@@
The PBW degeneration is compatible with the weighted ordering. ∎
@@ 3. Step 3
If there is $\alpha\in R^{+}$ with $u_{\alpha}\neq 0$, then $m\in U(\mathfrak{g}).e_{\alpha}\subseteq LT(M_{\lambda})$. If $u_{\alpha}=0$ for all $\alpha\in R^{+}$ and there exists $\alpha\in R^{+}$ with $t_{\alpha}\neq 0$. Using $h_{\alpha}-\lambda(h_{\alpha})\in M_{\lambda}$, showing that $LT(m)\in(LT(M_{\lambda}))$ reduces to showing that $LT(m^{\prime})\in(LT(M_{\lambda}))$ for all $m^{\prime}$ such that $LT(m^{\prime})\in LT(I(\lambda))\cap U(\mathfrak{n}^{-})$, that is $t_{\alpha}=u_{\alpha}=0$ for all $\alpha\in R^{+}$.

### 11 Jan 2024 v2

@@ Abstract.
We explicitly provide minimal Gröbner bases for simple, finite-dimensional modules of complex Lie algebras of types A and C, using a homogeneous ordering that is compatible with the PBW filtration on the universal enveloping algebras.
@@ 1. Introduction
as an element of $M_{\lambda}$. Since $f_{\alpha_{2}}^{2}$ is also contained, the set is not reduced. These computations are carried out using OSCAR [Osc] and the implementation of the Gröbner basis for other types can be found in the latter.
@@
Acknowledgments The first author gratefully acknowledges financial support by the DFG –Project-ID 286237555–TRR 195. The authors would like to express their gratitude to Thomas Breuer for his patient explanations of GAP.
@@ 2.3. Gröbner bases
For the following, we need to introduce a bit of notation: Let $\lambda=\sum m_{i}\omega_{i}\in P^{+}$, we set $\lambda_{i}=\sum_{j=i}^{n}m_{j}$. Then $\lambda$ is also the partition $(\lambda_{1}\geq\lambda_{2}\geq\ldots\geq\lambda_{n}\geq 0)$. Let $\mathbf{s}\in\mathbb{R}^{|R^{+}|}$ be supported on the roots of the subdiagram induced from $\alpha_{\ell},\ldots,\alpha_{k}$.
Since $I_{\lambda}$ is finitely generated and $U(\mathfrak{g})$ is a G-algebra (also known as PBW algebra or algebra of solvable type), a (left-)Gröbner basis exists for each adequate monomial ordering (see [KRW90]). We recall the monomial ordering on $U(\mathfrak{n}^{-})$ introduced in [FFL17] here:
A refinement of the natural ordering on positive roots is called a good ordering, for example $\alpha_{1,3}>\alpha_{1,2}>\alpha_{2,3}>\alpha_{1}>\alpha_{2}>\alpha_{3}$ for $\mathfrak{sl}_{4}$. We consider the induced degree reverse lexicographic ordering on $U(\mathfrak{n}^{-})$ (all root vectors have degree $1$ and $(1,0,0,1)<(0,1,1,0)$) and extend this to $U(\mathfrak{g})$ by setting $\mathfrak{n}^{-}>\mathfrak{h}>\mathfrak{n}^{+}$.
For type $A$ we set
@@
$\partial_{\mathbf{s}}\;-:=[e_{\ell,\ell}^{s_{\ell+1,\bullet}},[\;\cdots,[e_{\ell,k-1}^{s_{k,\bullet}},[e_{k,k}^{s_{\bullet,k-1}},[\;\cdots,[e_{\ell+1,k}^{s_{\bullet,\ell}},-]\cdots]]]\cdots]].$
To ease the notation, we do not recall $\partial_{\mathbf{s}}$ in type $C$ but refer to [FFoL11a, Theorem 3.4].
Let $\mathbf{s}$ be supported on a Dyck path $\mathbf{p}\in D_{i,j}$, then it is shown in [FFoL11, FFoL11a], that $LM(\partial_{\mathbf{s}}f_{i,j}^{\text{deg }\mathbf{s}})=f^{\mathbf{s}}$.
@@
The PBW degeneration is compatible with the monomial ordering. ∎
@@ 3. Step 3
If there is $\alpha\in R^{+}$ with $u_{\alpha}\neq 0$, then $m\in U(\mathfrak{g}).e_{\alpha}\subseteq(LT(M_{\lambda}))$. If $u_{\alpha}=0$ for all $\alpha\in R^{+}$ and there exists $\alpha\in R^{+}$ with $t_{\alpha}\neq 0$. Using $h_{\alpha}-\lambda(h_{\alpha})\in M_{\lambda}$, showing that $LT(m)\in(LT(M_{\lambda}))$ reduces to showing that $LT(m^{\prime})\in(LT(M_{\lambda}))$ for all $m^{\prime}$ such that $LT(m^{\prime})\in LT(I(\lambda))\cap U(\mathfrak{n}^{-})$, that is $t_{\alpha}=u_{\alpha}=0$ for all $\alpha\in R^{+}$.